3. A 10.0 meter length of straight conducting wire carries a current of 15.0 A. a) Find the resulting magnetic field at a distance of 10.0 cm from the wire. b) If this wire is now wrapped to form a circular coil with a 10.0 cm radius, find the value of the resulting magnetic field at the center of the coil. (Note that you will need to compute the number of loops N that can be made from a 10.0m length of wire; C = 2tr)
Answer #1

here,

length , l = 10 m

current , I = 15 A

a)

the resulting magnetic feild at a = 10 cm = 0.1 m , B = u0 * I /( 2 * pi * a)

B = 1.26 * 10^-6 * 15 /( 2 * pi * 0.1) T

B = 3 * 10^-6 T

b)

the number of turns , N = l/(2 * pi * a)

N = 10 /( 2 * pi * 0.1) = 16 turns

the value of the resulting magnetic feild , B = N * u0 * I /( 2 * a)

B = 16 * 1.26 * 10^-6 * 15 /( 2 * 0.1)

B = 1.5 * 10^-3 T